OpenStudy (anonymous):

write an equation in point-slope form for the line through the given point and the given slope (8,3); M=5

OpenStudy (anonymous):

In order to solve we need to find the formula for pont-slope form

OpenStudy (anonymous):

Point-Slope Formula: y – y1 = m(x – x1)

OpenStudy (anonymous):

x1=8 and y1=3 and m=5

OpenStudy (anonymous):

now just plug it all in!

OpenStudy (anonymous):

y-3=5(x-8)

OpenStudy (anonymous):

Use distributive property next so y-3=5x-40 Your going to multiply everything with 5 in the parenthesis

OpenStudy (anonymous):

Now you add +3 to both sides in order to cancel out the 3 on the left with Y. Because our goal is to keep Y by itself on the other side of the equation.

OpenStudy (anonymous):

y=5x-37 would be your answer :) hope I helped